# QPanda3: A High-Performance Software-Hardware Collaborative Framework for Large-Scale Quantum-Classical Computing Integration

### v1

@@ III-C Enhanced Performance
QPanda3 demonstrates significant performance improvements, with notable advantages primarily manifesting in efficient quantum program transmission and high-performance quantum circuit compilation.
@@
QPanda3 has designed a set of efficient protocols specifically for quantum program transmission, aiming to optimize the process of transmitting quantum programs between terminal and cloud devices, as well as among other devices. By standardizing the representation format and transmission process of quantum programs, this protocol significantly enhances the efficiency of users submitting tasks to quantum computing devices. Specifically, the protocol supports encapsulating and transmitting quantum programs in a compact intermediate representation form, avoiding the inefficiencies caused by non-uniform formats or excessive redundant information in traditional transmission methods. Additionally, the protocol employs efficient compression and encoding techniques, effectively reducing the amount of data during transmission, thereby saving communication bandwidth between devices and accelerating interaction speed. In practical applications, this protocol is not only suitable for task submission between user terminals and cloud-based quantum computing devices but also for program sharing and collaborative computing among different quantum computing devices. For example, in distributed quantum computing scenarios, multiple quantum devices can quickly exchange quantum programs and data through this protocol, achieving more efficient parallel computing and resource utilization. Furthermore, the design of this protocol takes into account the compatibility and scalability of quantum programs, adapting to the evolution needs of future quantum hardware and algorithms.
(2) Efficient Quantum Circuit Optimization and Compilation
The implementation and optimization of quantum algorithms still face numerous challenges, particularly in the construction and compilation of quantum circuits. As an advanced quantum computing framework, QPanda3 is dedicated to addressing these issues and providing efficient quantum circuit compilation tools. Through optimized implementation, QPanda3 significantly accelerates the process of constructing quantum logic circuits from Python code, enhancing the development efficiency of quantum algorithms. Secondly, the framework introduces the Transpiler lass, which enables device-oriented quantum circuit compilation. It dynamically adjusts compilation strategies based on the characteristics of the target hardware, thereby improving the execution efficiency and portability of quantum programs. Furthermore, QPanda3 optimizes the quantum circuit compilation process from multiple aspects, significantly enhancing the efficiency of quantum circuit compilation. With efficient quantum circuit compilation technology, QPanda3 provides strong support for the implementation and optimization of quantum algorithms.
III-D Software-Hardware Co-Design Oriented
QPanda3 is a modern quantum computing software development kit oriented towards software-hardware co-design. This is manifested in various key scenarios including quantum program design, quantum program transmission, quantum circuit compilation, quantum program analysis, and quantum program execution. As illustrated in Figure 3, QPanda3, running on both terminal and cloud devices, fully meets the critical requirements of quantum computing based on software-hardware co-design.
Figure 3: Workflow
(1) Cloud Execution and Inter-device Transmission of Quantum Programs
Due to the unique characteristics of quantum computers, the paradigm of terminal design for quantum programming coupled with cloud execution has been widely accepted and has found numerous practical applications. A hallmark of this paradigm is the separation of devices used for program design from those used for program execution, with the two interconnected and collaborating through communication links. Quantum programs are transmitted from the design end to the computation end via these communication links. QPanda3 is not merely a tool for quantum programming; rather, it is a collaborative tool that adheres to the aforementioned paradigm. On one hand, QPanda3 provides convenient access to the quantum cloud, allowing quantum programs designed on QPanda3 to be directly executed on various computing devices within the quantum cloud. On the other hand, to optimize the efficiency of quantum program transmission between devices, QPanda3 has designed a compact intermediate representation for quantum programs. Additionally, QPanda3 offers tools for converting between quantum programs expressed in high-level programming languages and this intermediate representation. Furthermore, serving as an interactive interface for users to access quantum computing devices, QPanda3 enables users to obtain relevant information about the computing devices.
(2) Hardware-Software Co-Design for Optimized Compilation of Quantum Circuits
Quantum circuit compilation is the process of transforming quantum logic circuits into sequences of machine instructions. This process, which is a software-based procedure, necessitates a comprehensive consideration of the actual connection topology of physical qubits and the physical operations that the device employs to control these qubits, with such operations typically corresponding to a limited set of quantum gates. Clearly, quantum circuit compilation represents a typical scenario of hardware-software co-design. The extent of various optimizations applied during the quantum circuit compilation process influences the degree to which quantum programs utilize device performance, and directly impacts the correctness of quantum computation results. QPanda3 decouples key steps in quantum circuit compilation, including qubit mapping, qubit routing, and machine instruction generation, thereby enabling users to compile and generate machine instruction sequences that can be efficiently executed on specified devices, based on different physical qubit layouts and native quantum gate sets. During the circuit compilation process, QPanda3 implements multiple software-level optimization strategies, some of which accelerate compilation speed, while others enhance the execution performance of the compiled circuits.
(3) Hardware-Software Co-Design for Quantum Program Analysis and Design
The goal of quantum programming is to develop quantum programs that can solve target problems, fully utilize computational resources, and exhibit acceptable quantum computing performance. For the quantum computing component, the key to quantum programming lies in obtaining efficient quantum logic circuits.
QPanda3 enables users to design quantum programs using a high-level programming language (Python) and a user-friendly intermediate representation (OriginIR). QPanda3 provides a variety of quantum simulators, which can be deployed on the quantum cloud as services for users. These simulators are also built into QPanda3, allowing users to run them using the computing resources of their personal computers (CPUs, GPUs). Quantum simulators on the cloud and on personal computers can be used to verify whether the theoretical execution results of quantum programs match expectations, thereby guiding the further design of quantum programs. Collaborating with local or cloud computing resources to validate the correctness of quantum programs represents a classic hardware-software co-design scenario. There may be multiple implementations of quantum programs that solve the same problem, and different quantum programs may exhibit varying performance on the same device. It is of great significance to select an appropriate implementation (or subroutine) of a quantum program that is tailored to the target problem and the available computing device. Such task- and device-adapted quantum programs can obtain more accurate results more efficiently, aligning more closely with the goals of quantum programming. This type of quantum program design relies on device-based quantum program analysis. QPanda3 provides this analysis capability, facilitating users in designing more efficient quantum programs. Specifically, QPanda3 offers quantum program analysis tools that are oriented towards quantum computing performance.
IV Representation and Transmission of Quantum Programs
IV-A User-Friendly Intermediate Representation – OriginIR
@@
IV-A1 Compatibility with QPanda2
@@
IV-A2 Geared Towards Researchers
@@
IV-B Intermediate Representation for Quantum Program Transmission - OriginBIS

### v2

@@ III-D Enhanced Performance
QPanda3 delivers notable performance enhancements, especially in quantum program transmission and high-efficiency circuit compilation.
@@
To address the data representation and communication bottlenecks inherent in large-scale and distributed quantum–classical hybrid computing workflows, QPanda3 introduces a unified, high-efficiency quantum program transmission protocol. This protocol enables robust interaction between user clients, cloud-based orchestration layers, and quantum backends.
At its core lies the Unified Quantum Program Representation Module, which abstracts quantum instructions and metadata into a compact, binary-encoded intermediate representation (IR). This structure-preserving design facilitates compatibility across heterogeneous hardware and execution environments, while also enabling fast serialization and deserialization.
Figure 4: Architecture of the compilation module with lightweight data structures and module-level transformations.
As illustrated in Figure 4, the IR undergoes multi-level compression, including gate dictionary encoding, instruction fusion at the bitstream level, and entropy-aware payload optimization. These strategies collectively reduce the size of the transmitted data by up to 80% compared to text-based formats such as OpenQASM, thereby enhancing throughput and minimizing latency in cloud-device quantum computing pipelines.
Optional encryption and integrity verification mechanisms—via hashing and secure payload encapsulation—are incorporated to safeguard against tampering in multi-tenant or federated quantum systems.
(2) High-Performance Quantum Circuit Compilation
A customized data structure design method tailored to the characteristics of the quantum compilation workflow is proposed to address the issue of redundant intermediate information management in instruction reordering, topology-constrained mapping, and gate fusion optimization. The core of this approach lies in the construction of the quantum circuit’s directed acyclic graph (DAG), where lightweight operation units and contextual annotation mechanisms are introduced. Logical gate operations are organized into pipelined fragments with localized context, enabling structured phase buffers and operation dependency chains to support efficient access and dynamic updates of critical dependencies throughout gate-level optimization, scheduling, and topology mapping. This method significantly enhances compilation speed while preserving compilation quality, thereby meeting the rapid deployment requirements for large-scale quantum systems with 1000+ qubits.
Figure 5: Architecture of the compilation module with lightweight data structures and module-level transformations.
The Figure 5 introduces a contextual annotation mechanism that segments the circuit into pipelined fragments with localized context. Each segment is annotated with metadata such as segment identifiers, topological constraints, and temporal execution windows, enabling fine-grained control over optimization boundaries. These annotated fragments are managed within a stage buffer, which acts as an optimization context pool for subsequent compilation stages.
A dependency-linked list structure is maintained to support cross-segment dependency tracking. This linked representation allows for low-latency traversal and update of gate dependencies during optimization. The integrated design of the DAG, stage buffer, and dependency tracking supports structured construction of operation dependency chains, which facilitates efficient updates and localized scheduling decisions throughout various optimization passes.
The optimization layer leverages this structured data representation to enable a sequence of specialized transformations, including:
  • Pipeline-stage layout optimization, which improves spatial coherence of gate execution,
  • Parallel gate fusion, which merges compatible gates across segments to reduce circuit depth,
  • Constraint-driven topology mapping, which ensures hardware compliance, and
  • Dependency-aware gate translation, which refines gate translating under strict causal constraints.
This architecture demonstrates significant improvements in compilation throughput, supporting compilation for systems with 1000+ qubits. By integrating operation structure, temporal context, and dependency linkage into a unified data-driven framework, this approach not only accelerates the compilation pipeline but also ensures high-quality output circuits that are compliant with the constraints of realistic NISQ hardware.
III-D1 Gate Fusion
Gate fusion is performed in a two-stage iterative pipeline. First, adjacent two-qubit gates are fused in parallel across pipeline segments to form localized composite gate blocks. Each fused block is then analyzed for potential decomposition into a tensor product of single-qubit gates. If such decomposition is feasible, the resulting single-qubit gates are merged into the surrounding context, enabling further compression of circuit depth.
This fusion–decomposition–merging loop, driven by explicit dependency tracking and localized segment annotations, yields a structurally optimized two-qubit gate layer that significantly reduces overall circuit depth.
III-D2 Layout Optimization
Layout optimization begins with a subgraph reconstruction algorithm that attempts to map logical circuit substructures onto hardware topologies within a bounded runtime. This mapping process considers spatial gate alignment, dependency constraints, and SWAP minimization.
If an exact subgraph match is not found within the allowed time frame, the compiler switches to a noise-aware mapping strategy. In this fallback mode, qubit blocks are selected based on backend calibration data, prioritizing hardware regions with higher gate fidelity to improve the robustness and execution fidelity of the final circuit.
III-D3 Topology Mapping
Topology mapping is built upon the LightSabre algorithm, which offers superior scalability for deep circuits and large hardware graphs. To improve mapping efficiency, a lightweight data structure is designed to represent qubit connectivity and cost metrics.
This structure maintains sparse neighborhood graphs augmented with local path caches and dynamic edge costs, allowing fast constraint evaluation and route selection without the overhead of global matrix traversal. The result is a high-throughput mapping process that scales effectively to devices with 1000+ qubits.
III-D4 Gate Translation
Gate translation enables backend-agnostic circuit generation by converting logical gate sets into native hardware instructions. A graph-based dependency model is used to capture gate relationships, allowing flexible decomposition of complex operations into target-compatible primitives.
To mitigate the depth overhead introduced by redundant single-qubit gates—often a side effect of two-qubit gate decomposition—the compiler adopts a fused translation–optimization strategy, where gate synthesis and simplification are co-executed in a single pass.
For large-scale quantum programs, this process is further accelerated via parallelism, allowing multiple gate segments to be translated and optimized concurrently.
(a) OriginIRs’ Encode Time
(b) OriginIRs’ Decode Time
(c) OriginIRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) OriginBIS’ Post-encoding Size
Figure 6: For Circuit Count
IV Experiments
IV-A Performance Experiments of OriginBIS
(a) OriginIRs’ Encode Time
(b) OriginIRs’ Decode Time
(c) OriginIRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) OriginBIS’ Post-encoding Size
Figure 7: For Circuit Depth
(a) OriginIRs’ Encode Time
(b) OriginIRs’ Decode Time
(c) OriginIRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) OriginBIS’ Post-encoding Size
Figure 8: For Qubit Count
IV-A1 Evaluation Metrics
(1) Encode Time and Decode Time
(2) Post-encoding Size
IV-A2 Experimental Setup
TABLE I: Experimental Setup For Quantum Program Transmission
Experiment Circuit Count Circuit Depth Qubit Count
For Circuit Count Independent Variable 500 72
For Circuit Depth 500 Independent Variable 72
For Qubit Count 500 500 Independent Variable
IV-A3 Experimental Results and Analysis
(1) For Circuit Count
(2) For Circuit Depth
(a) OriginIRs’ Encode Time
(b) OriginIRs’ Decode Time
(c) OriginIRs’ Post-encoding Size
(d) OriginBIS’ Encode Time
(e) OriginBIS’ Decode Time
(f) OriginBIS’ Post-encoding Size
Figure 9: For Gate Count
(3) For Qubit Count
(4) Gate Count
Figure 10: Gate Count for Circuit Count, Circuit Depth and QBit Count
Figure 11: Experimental Results About Circuit Compilation
IV-B Compilation Efficiency Comparison Experiment
IV-B1 Experiment Setup
TABLE II: SDKs
SDK Version
amazon-braket-sdk(braket) 1.91.0
bqskit 1.2.0
cirq 1.4.1
pytket(tket) 2.0.1
qiskit 2.0.1
qiskit_ibm_transpiler 0.39.0
pystaq 3.5
pyqpanda3 0.3.0
TABLE III: Software and Hardware Environments
Type Info
CPU Intel(R) Xeon(R) Platinum 8336C CPU @ 2.30GHz
Memory 2.0TB
Operating System Ubuntu 20.04.5 LTS
Python 3.11.11
IV-B2 Evaluation Metrics
IV-B3 Experimental Results And Analysis
SDK Passed Failed XFailed Skipped Total
braket 7 2 0 1057 1066
cirq 10 2 0 1054 1066
qiskit 1044 0 0 22 1066
QPanda3 1016 28 0 22 1066
The performance of various SDKs in circuit construction is illustrated in Figure 12. It can be observed that QPanda3, qiskit, and cirq have successfully passed all seven test cases. For the four test cases, namely test_multi_control_circuit, test_bigint_qasm2_import, test_param_circSU2_100_build, and test_clifford_build, QPanda3 demonstrated the shortest execution time. In the case of the three test cases test_QV100_qasm2_import, test_param_circSU2_100_build, and test_QV100_buildQPanda3 took longer than qiskit but significantly less time compared to cirq, braket, and bqskit. For the test case test_DTC100_set_build, QPanda3’s execution time was shorter than that of qiskit but longer than cirq’s, ranking as the second shortest. This figure highlights the advantages of QPanda3 in circuit construction.
The Benchpress literature indicates that Manipulate is a crucial metric for evaluating SDKs. As illustrated in Figure 13, only QPanda3 and qiskit have successfully passed all four test cases. In the test_QV100_basis_change test case, QPanda3 demonstrated the shortest execution time. For the three test cases: test_random_clifford_decompose, test_multi_control_decompose, and test_DTC100_twirling, QPanda3’s execution time slightly exceeded that of qiskit, with a minimal difference. This figure suggests that QPanda3 holds an advantage in terms of the Manipulate metric.
IV-C Performance Comparison Experiment for Gradient Computation in Variational Quantum Circuits
Analysis of Gradient Computation Performance and Circuit Depth in 12-Qubit Systems
Figure 14: Circuit Depth Impact (Fixed 12 Qubits)
The performance analysis of gradient computation in 12-qubit systems corresponding to Figure 14 demonstrates that quantum framework selection critically determines computational efficiency. PyQPanda3 exhibits significant time advantages, requiring only 0.0088 seconds average time at 2-layer circuit depth compared to Qiskit’s 0.232 seconds - a 26-fold difference. When depth increases to 6 layers, PyQPanda3’s time merely grows to 0.0241 seconds (2.7x increase), substantially lower than Qiskit’s 0.6097 seconds (2.6x) and DeepQuantum’s 0.1680 seconds (2.9x). Notably, frameworks maintain a consistent performance hierarchy: PyQPanda3 consistently delivers optimal performance, with computation times 40%-50% lower than second-ranked PennyLane-Lightning (0.0241s vs 0.0362s at 6 layers), while Qiskit persistently trails. During depth scaling, PyQPanda3 shows the smallest absolute time increment (+0.0153s), equivalent to just 4% of Qiskit’s increment (+0.3777s), enabling 97.6% time savings in deep circuits. Experimental data confirms PyQPanda3’s superior depth-scaling stability, with significantly lower time-complexity growth rates than comparable frameworks, providing an effective computational efficiency solution for deep quantum circuits.
Comparative Analysis for Fixed Qubit Count and Circuit Depth
Figure 15: Framework Comparison 8qubits 4layers
Figure 16: Framework Comparison 16qubits 6layers
Based on the comparative analysis of Figure 15 and Figure 16, a significant and stable statistical pattern emerges in gradient computation times across quantum frameworks. Under 8-qubit/4-layer conditions, PyQPanda3 demonstrates exceptional performance (0.0031s), requiring only 21.7% of the time needed by second-ranked PennyLane-Lightning (0.0143s), while Qiskit’s duration (0.3090s) reaches 99.7x that of PyQPanda3. When scaling to 16-qubit/6-layer configurations, the performance ranking remains strictly consistent: PyQPanda3 (0.1116s) maintains the lowest time, outperforming PennyLane-Lightning (0.2134s) by 47.7%, with Qiskit (2.0085s) requiring 18x more time than PyQPanda3.
Notably, the performance differences exhibit systematic hierarchical characteristics: PyQPanda3 and PennyLane-Lightning consistently form an efficient cluster, collectively accounting for 16.1% and 7.5% of total computation time under the two respective scales; DeepQuantum and PennyLane-Default comprise a mid-tier cluster; while Qiskit independently constitutes an inefficient unit, dominating 82.9% (8-qubit) and 68.2% (16-qubit) of total time. As computational scale increases, PyQPanda3’s absolute advantage over Qiskit expands from 0.3059s to 1.8969s, though its relative advantage multiplier decreases from 99x to 18x, indicating nonlinear evolution of performance differentials with scaling. These statistical patterns confirm that quantum framework performance disparities maintain intrinsic stability regardless of computational scale.
Time Analysis of Gradient Computation for Fixed 4-Layer Quantum Circuits
Figure 17: Qubit Scalability 4layers
Analysis of gradient computation time statistics for fixed 4-layer quantum circuits reveals the following patterns: As qubit count increases from 4 to 16, all frameworks exhibit exponential time growth (manifesting as linear trends in logarithmic coordinates), yet with significantly different growth slopes. PyQPanda3 demonstrates the mildest growth (approximately 100x increase from 4 to 16 qubits), while Qiskit shows the steepest escalation (approximately 10x increase). PennyLane-Default and DeepQuantum exhibit intermediate growth slopes with stable differentials, maintaining a consistent time ratio of $\approx$1:1.2 throughout the scaling process.
The performance differentials between frameworks exhibit dual-scale dependency: Regarding absolute differences, the maximum time gap continuously widens with increasing qubits (from 0.099s between PyQPanda3 and Qiskit at 4 qubits to 0.9s at 16 qubits); whereas in relative terms, the advantage of efficient frameworks diminishes with scale (the PyQPanda3/Qiskit time ratio decreases from 1:100 at 4 qubits to 1:10 at 16 qubits).
Comparison with PennyLane-Lightning
Figure 18: Speedup Heatmap vs Pennylane-lightning
Based on the analysis of the speedup heatmap 18, QPanda3 exhibits significant systematic performance advantages over Pennylane Lightning. In shallow small-scale quantum circuits (depth $\leq$6 layers, qubits $\leq$8), QPanda3 demonstrates prominent acceleration effects, reaching up to 26.0x (2-layer/4-qubit configuration). However, this advantage exhibits dual-decay characteristics with parameter variations: when fixing qubit count, the speedup ratio decreases sharply with increasing depth (e.g., for 4-qubit circuits: from 26.0 at 2 layers to 1.2 at 12 layers); when fixing depth, the speedup ratio decays progressively with qubit scaling (e.g., at 2-layer depth: from 26.0 at 4 qubits to 6.5 at 16 qubits).
The impact of circuit depth on acceleration effects is particularly significant: a depth of 8 layers constitutes a critical turning point, where speedup ratios are generally $>$4.0 below this threshold (e.g., 5.3 for 6-layer/4-qubit circuits) but typically $<$2.0 above it (e.g., only 1.6 for 10-layer/4-qubit circuits). A qubit count of 12 emerges as another demarcation point-below this scale, speedup ratios exhibit drastic fluctuations (maximum 73% decrease from 4 to 8 qubits), while above this scale they tend to stabilize ($<$5% variation between 14-16 qubits).
The spatial distribution of performance advantages exhibits a clear hierarchical structure: strong acceleration zones ($>$5.0) are concentrated in the upper-left triangular region (depth $\leq$6 layers & qubits $\leq$8), weak acceleration zones (1.2-2.0) occupy the lower-right triangular region (depth $\geq$10 layers or qubits $\geq$14), and transition zones (1.6-4.9) are distributed along the diagonal. This distribution pattern indicates that the inhibitory effect of circuit depth on acceleration surpasses that of qubit scale, converging to a 1.2-2.0x performance gap between the two frameworks when depth $\geq$10 layers or qubits $\geq$14.
Comparison with Qiskit
Figure 19: Speedup Heatmap vs Qiskit
Analysis based on heatmap 19 data indicates that QPanda3 demonstrates significant performance advantages compared to Qiskit, with its acceleration effects exhibiting clear parameter-dependent patterns. In medium-depth small-scale quantum circuits (48 layers depth, 610 qubits), QPanda3 achieves super-strong acceleration, peaking at 157.3 times (6 layers depth/6 qubits). This advantage shows asymmetric attenuation with circuit parameter changes: the attenuation intensity caused by increasing depth (up to 83%) is significantly higher than the impact of increasing qubit count (up to 45%). Notably, a depth of 8 layers constitutes a critical threshold-below this value, the speedup ratio is generally $>$68.6 times (e.g., 99.2 times for 6 layers depth/8 qubits), while above this value it sharply decreases to $<$27.8 times (e.g., 24.6 times for 10 layers depth/4 qubits).
The impact of qubit scale on acceleration exhibits phased characteristics: when the number of qubits $\leq$10, the speedup ratio fluctuates drastically with increasing depth (e.g., from 70.4 to 24.6 for 4-qubit circuits when depth increases from 2 to 10 layers); whereas when qubits $\geq$12, the speedup ratio shows blunted responses to parameter variations (e.g., 16-qubit circuits maintain only 17.118.0x acceleration across different depths). This differential response forms three characteristic spatial distribution zones: super-strong acceleration zone ($>$80x) concentrates in the rectangular region with depth $\leq$6 layers and qubits $\leq$10; strong acceleration zone (40-99x) extends to depth 8 layers / qubits 14; stable acceleration zone (17-28x) covers the lower-right triangular region where depth $\geq$10 layers or qubits $\geq$12.
Even under deep-layer large-qubit-scale conditions (16 layers depth/16 qubits), QPanda3 maintains stable acceleration advantages exceeding 17x, with fluctuation amplitude $<$20%. This phenomenon indicates that while circuit depth exerts stronger inhibitory effects on acceleration than qubit scale, QPanda3 consistently demonstrates significant performance advantages across the entire parameter space, particularly exhibiting groundbreaking acceleration capabilities in shallow-to-medium quantum circuits.
V Conclusion
This paper focuses on efficient compilation and execution in quantum computing, introducing OriginIR and OriginBIS as intermediate representations to enhance programmability, transmission efficiency, and execution performance. Experimental results demonstrate that OriginBIS significantly outperforms OpenQASM 2.0 in encoding and decoding speed as well as information capacity. Additionally, QPanda3 surpasses Qiskit in quantum circuit construction, operation execution, and compilation speed, exhibiting exceptional performance improvements, particularly in large-scale quantum circuit compilation tasks. QPanda3 provides fundamental and efficient functionalities related to variational quantum circuits, offering robust support for quantum-classical hybrid computation. These advancements lay a solid foundation for the engineering applications of future quantum computing.
VI Future Work
Appendix A Representation and Transmission of Quantum Programs
A-A User-Friendly Intermediate Representation-OriginIR
@@
A-A1 Compatibility with QPanda2
@@
A-A2 Geared Towards Researchers
@@
A-B Intermediate Representation for Quantum Program Transmission - OriginBIS